# CoA: Chain-of-Action for Generative Semantic Labels

### v1

@@ 4.2 Quantitative results
An interesting observation from the table is that overall, VQA-based approaches perform less effectively than caption-based methods in $\text{M}_{\text{clip}}$. This is primarily because the GSLs task emphasizes identifying all entities relevant to an image. Single-question VQA interactions often lead LMMs to focus on specific local details within an image, whereas captioning guides the model to capture a more comprehensive set of features, resulting in outputs that encompass all relevant entities. This insight influenced the design of our action chain, where the first action is dedicated to obtaining a broad image description. Our experiments validate the effectiveness of this setup.
4.3 Effectiveness analysis of CoA
To validate the effectiveness of the CoA, we decomposed it into multiple independent components. Since Action 5 serves as the primary step for the GSLs task, responsible for identifying all entity labels in the test image, we used it as the baseline. We then incrementally added each subsequent component to the CoA and evaluated the performance impact after each addition. The results of these evaluations are shown in Table 4. From Table 4, it is evident that each added component contributes to significant improvements in both $\text{M}_{\text{clip}}$ and $\text{M}_{\text{ram}}$ metrics. Notably, the complete action chain outperforms using only Action 5, achieving an increase of $+13.46\%$ in the $\text{M}_{\text{clip}}$ metric and $+2.15\%$ in the $\text{M}_{\text{ram}}$ metric. These findings strongly highlight the performance enhancements brought by the CoA, which proves the effectiveness of the proposed method.
4.4 Applicability for various LMMs
To evaluate the adaptability of CoA across various LMMs, we conducted experiments by changing the base model and testing its performance on different LMMs. Specifically, we employed LLaVA-1.5-7B and LLaVA-1.5-13B as base model. Although LLaVA-1.5-13B is generally more capable than LLaVA-1.5-7B, integrating CoA with LLaVA-1.5-13B still resulted in significant improvements. Notably, there was an increase of $+14.28\%$ in the $\text{M}_{\text{clip}}$ metric and $+12.1\%$ in the $\text{M}_{\text{ram}}$ metric. These findings demonstrate the adaptability and effectiveness of the proposed CoA across diverse LMMs, highlighting its capability to integrate seamlessly with different model architectures.
4.5 Additional ablations study
Multiple interaction vs single interaction. To validate the advantage of multiple interactions, we designed an experiment inspired by the concept of In-Context Learning. This involved merging the multiple actions of CoA into a single interaction (i.e., guiding LMMs to simultaneously consider captions, the appearance of entities, and relationships between entities within a single interaction; the detailed merged template is provided in Appendix F). The comparison of single-interaction performance versus multi-interaction CoA is reported in Table 6. The results clearly demonstrate that multiple interactions provide significant performance benefits over a single-step approach, attributed to the more granular processing and well-defined task objectives that facilitate more accurate outputs from LMMs.
Time cost of CoA. Given that the proposed CoA method involves multiple interactions with LMMs, it is essential to evaluate the time cost associated with the overall execution process. We analyze the computational efficiency of CoA versus VQA-based LMMs and Caption-based LMMs and report their inference time in Table 7. Notably, our approach maintains a comparable execution efficiency despite multiple interactions. This is attributed to the simplicity of each individual interaction, ensuring that the total inference time remains within an acceptable range.
@@ 4.5 Additional ablations study
      VQA       BLIP-2 [24]       6650 $\pm$ 117
      LLaVA-1.5 [29]       753 $\pm$ 14
      MiniGPT-4 [58]       11216 $\pm$ 335
      Caption       BLIP-2 [24]       6870 $\pm$ 177
      LLaVA-1.5 [29]       3073 $\pm$ 152
      MiniGPT-4 [58]       14760 $\pm$ 1433
      CoA (Our)       6110 $\pm$ 149
Table 7: Computational cost of different methods on the Split-3 of VOC dataset.
With / Without Ram filtering. During the experimental process, we observed that beyond using the RAM model for designing evaluation metrics, its confidence scores could also be leveraged to filter generated targets. To validate this, we conducted experiments comparing results with and without RAM-based filtering, as reported in Table 8. The results demonstrate that applying this filtering method significantly improves the accuracy of the GSLs task.
Dataset Method Split-0 Split-1 Split-2 Split-3 Avg
VOC CoA 75.18 78.54 79.69 76.54 77.49
CoA w/ filter 88.19 88.84 89.64 88.25 88.73
COCO CoA 82.04 80.37 83.02 76.51 80.49
CoA w/ filter 91.88 91.03 93.79 88.32 91.11
NUS CoA 70.43 70.99 70.23 68.57 70.06
CoA w/ filter 84.94 84.93 84.39 83.67 84.46
Table 8: Comparison of CoA without / with RAM filtering in $\text{M}_{\text{ram}}$ metric on VOC dataset.
5 Conclusion
@@
In this work, unlike traditional VLMs that rely on predefined label sets, we propose a new task, GSLs, which generate comprehensive and contextually relevant labels from an unconstrained semantic space. Besides, we introduced an innovative Chain-of-Action (CoA) method designed to handle the challenging GSLs task. By breaking down the generative task into sequential actions that iteratively extract and refine key information, our method ensures that LMMs generate more accurate and complete semantic labels. Finally, extensive evaluations on widely-used benchmark datasets validate the efficacy of CoA, showcasing significant performance gains across multiple key metrics. This highlights the potential of our method to improve label generation in dynamic, evolving environments.
Appendix A Instruction templates of VQA LMMs
In this section, we present several instruction templates used within the VQA-based LMMs baselines, as illustrated in Fig. 5. For our study, these templates were employed to directly obtain semantic labels present in the images when applying the VQA-based baseline methods. We evaluated the effectiveness of these different instruction templates and selected the optimal configuration for comparison. The results of these comparisons are reported in Table 1 and Table 2 in the main manuscript.
Figure 5: The instruction templates of VQA LMMs.
Appendix B Details of filtering strategy
In this section, we provide additional details on how to filter the initial list of semantic labels after performing the Caption Action. For simplicity and efficiency, we adopt a straightforward approach. Specifically, we first create a set of all words present in the generated caption. This set is then filtered step by step. Using the inflect package, we perform part-of-speech tagging to identify and retain only noun-type semantic labels. Finally, we remove irrelevant entries (e.g., “image,” “photo,” “logo,” etc.) that are unrelated to the GSLs task, and then apply deduplication and singular/plural normalization. It is worth noting that this filtering strategy is a basic, streamlined process, and it is applied consistently across all Caption-based LMMs baseline methods.
Appendix C Details for the models
In this section, we provide a detailed description of the various LMMs employed in the baseline methods. The specific details are outlined as follows.
  1. $\bullet$
    LLaVA-1.5. The LLaVA [29] architecture stands out as a state-of-the-art (SOTA) LMM method, notable for its effective integration of CLIP visual features and LLM language tokens through a straightforward linear projection into a shared embedding space. LLaVA is fine-tuned using the LLaVA-Instruct-158k dataset, which consists of images paired with conversational prompts, detailed descriptions, and complex reasoning responses, achieving superior visual alignment compared to basic image-text pairs. In this study, we evaluate LLaVA-1.5 [29], an enhanced version that offers improved baselines. Key advancements over the original model include: (1) replacing the linear projection with a multi-layer perceptron (MLP) and (2) pretraining on a more diverse set of datasets.
  2. $\bullet$
    BLIP-2. BLIP-2 [24] is an advanced vision-language model designed to bridge the gap between visual and linguistic representations by incorporating a two-stage learning framework. The model begins by using a lightweight visual encoder (such as CLIP [34]) to extract image features, followed by a query-based cross-modal transformer that effectively aligns these visual representations with text. This architecture allows BLIP-2 to perform zero-shot and few-shot learning tasks with enhanced efficiency and scalability. By leveraging pretraining strategies that integrate visual context with language models like OPT or GPT-style backbones, BLIP-2 achieves strong performance in both image-to-text and multimodal tasks. Its innovative design reduces the computational overhead of extensive vision-language alignment training, making it a competitive choice for tasks requiring detailed image understanding and nuanced text generation.
  3. $\bullet$
    InstructBLIP. InstructBLIP [9] is a cutting-edge vision-language model designed to align multimodal understanding with human-like instruction following. The architecture builds upon the foundation of BLIP-2 [24], integrating a vision encoder (e.g., CLIP [34]) with an enhanced cross-modal transformer that incorporates instruction-based pretraining. InstructBLIP distinguishes itself by being fine-tuned with a diverse range of multimodal instructions, enabling the model to better understand and respond to complex tasks that involve visual and textual cues. This enables more coherent and contextually relevant responses, aligning model outputs with user prompts across various scenarios, from question-answering and image captioning to reasoning and complex interactions. By incorporating instruction-tuned objectives, InstructBLIP advances performance in both zero-shot and fine-tuned benchmarks, demonstrating superior adaptability and alignment with user intent compared to standard vision-language models.
  4. $\bullet$
    MiniGPT-4. MiniGPT-4 [58] is a vision-language model designed to emulate the capabilities of large multimodal models, such as GPT-4 [1], while being more lightweight and efficient. It integrates a visual encoder, typically based on advanced vision transformers like CLIP [34], with a language model through a linear transformation layer, aligning visual embeddings with language tokens. MiniGPT-4 undergoes a two-stage training process: first, a pretraining phase using image-text pairs to establish a foundational understanding of multimodal data, and second, an instruction-tuning phase to refine its responses and better align them with human interactions. This instruction-tuned process enhances the model’s performance on tasks that involve image descriptions, visual question answering, and complex reasoning. MiniGPT-4 is notable for its reduced computational requirements while maintaining strong alignment and contextual comprehension capabilities, making it a practical alternative for applications requiring efficient multimodal reasoning.
Appendix D Evaluation metrics
Given that the GSC task requires identifying multiple entities, we aim for model predictions to be as comprehensive and accurate as possible. Therefore, we designed two evaluation metrics that separately assess these aspects. Specifically, leveraging the generalization capability of the CLIP model, we used the prompt $P_{com}:$ “This image contains […]’ to capture the comprehensiveness of predicted entities. Next, given the predicted entity list $Y_{i}$ and the manually annotated entity list $\bar{Y}_{i}$, we compute their respective similarities to the image. Then, we can obtain the comprehensive score $CS_{i}$ or the predicted labels of each test image $I_{i}$ from the following formula:
$\small CS_{i}=\mathbf{1}\left[cos\langle T(P_{com}(Y_{i})),V(I_{i})\rangle>cos\langle T(P_{com}(\bar{Y}_{i})),V(I_{i})\rangle\right],$ (1)
where $\mathbf{1}[\cdot]$ denotes the indicator function, $T(\cdot)$ and $V(\cdot)$ denotes the text and image encoder of CLIP model. A prediction is scored only if the similarity of the model-generated prompt exceeds that of the prompt based on manual annotations. Here, we utilize the label set provided by the dataset itself as the manual-annotated labels. Then, we can obtain the $M_{clip}=\sum_{i}^{N}CS_{i}/N$ to assess the coverage of predicted labels, where $N$ denotes the number of test images.
To evaluate the accuracy of each predicted label, we leverage the RAM model to score each predicted entity label. RAM, a large-scale multi-label model trained on extensive multi-label datasets, has demonstrated significant recognition accuracy in multi-label tasks. Given a test image and a set of labels, RAM calculates a confidence score for each label in the list, filtering out labels that align with the image based on confidence thresholds. The accuracy score $AS_{i}$ of the predicted label list is then computed using the following formula:
$\small AS_{i}=Sigmod(\sum_{j}^{L}H(RAM(Y_{i}^{j}))),$ (2)
where $RAM(Y_{i}^{j})$ denotes the confidence score calculated by RAM model for the label $j$ from predicted labels set $Y_{i}$, $L$ denotes the number of $Y_{i}$, and
$\small H(RAM(Y_{i}^{j})))=\left\{\begin{aligned} 1\quad if\quad RAM(Y_{i}^{j})\geqslant\sigma\\ -1\quad if\quad RAM(Y_{i}^{j})<\sigma\\ \end{aligned}\right.,$ (3)
where $\sigma$ denotes the threshold of confidence score. Following the RAM configuration, we set the confidence threshold to $0.73$. A straightforward scoring rule is applied: one point is added for each correct prediction, and one point is subtracted for each incorrect prediction. Finally, we calculate the $M_{ram}=\sum_{i}^{N}AS_{i}/N$ to assess the accuracy of the predicted labels.
Appendix E Details for the datasets
We employ widely used datasets for evaluation, including PASCAL VOC 2012 (VOC) [12], MS-COCO 2014 (COCO) [26], and NUS-WIDE (NUS) [7]. The more details of these datasets are outlined as follows.
  1. $\bullet$
    PASCAL VOC 2012 (VOC) [12]: The PASCAL VOC 2012 dataset is a benchmark for evaluating algorithms in visual object recognition and detection. It is part of the PASCAL Visual Object Classes Challenge, a series of competitions designed to push the boundaries of object detection and classification research. The dataset comprises 20 object categories spanning everyday scenes, such as vehicles, animals, and household items. It includes a training/validation set with over 11,000 annotated images and approximately 27,000 object instances, providing a robust foundation for developing and benchmarking algorithms. Each image in the dataset is annotated with high-quality, pixel-wise segmentation masks and object bounding boxes, enabling multi-task learning in both image classification and semantic segmentation. VOC remains widely used due to its diverse visual content, standardized evaluation metrics, and significant influence on advancing deep learning methods in computer vision. Source: http://host.robots.ox.ac.uk/pascal/VOC/
  2. $\bullet$
    MS-COCO 2014 (COCO) [26]: The MS-COCO 2014 dataset is a widely recognized benchmark in the computer vision community, designed to facilitate research in object detection, segmentation, and captioning. It contains over 300,000 images with rich contextual information, depicting 80 object categories commonly found in everyday scenes. The dataset is meticulously annotated, featuring more than 2.5 million labeled instances that include precise object boundaries and segmentation masks, enabling comprehensive multi-task learning. In addition to detection and segmentation tasks, COCO provides a unique image captioning component, where each image is paired with multiple human-generated descriptions to support advancements in image-to-text models. Its diverse and challenging set of images, captured in realistic and complex environments, has made COCO a cornerstone for training and evaluating modern deep learning algorithms, significantly influencing state-of-the-art performance in visual understanding tasks. Source: http://images.cocodataset.org/zips/val2014.zip
  3. $\bullet$
    NUS-WIDE (NUS) [7]: The NUS-WIDE dataset is an extensive benchmark designed to facilitate research in large-scale image classification, multi-label annotation, and retrieval. NUS comprises over 260,000 images sourced from Flickr, annotated with 81 concept labels spanning various categories such as objects, scenes, and events. Each image in the dataset is tagged with one or more labels, supporting multi-label learning and evaluation. The annotations are derived using both automated and manual verification processes, ensuring label reliability and diversity. NUS also includes pre-extracted low-level visual features, providing a foundation for baseline experiments. This dataset is notable for its real-world distribution, reflecting the challenges of noisy data and varying image quality found in online media. Its breadth and comprehensive labeling make it a valuable resource for advancing multi-label classification, image retrieval, and large-scale visual recognition research.
Appendix F Merged template for single interaction
To thoroughly validate the advantages offered by multiple interactions, we designed an experiment based on the principles of In-Context Learning [15, 32]. In this experiment, the sequential steps of the proposed CoA framework were combined into a single interaction. This approach guided the LMMs to concurrently process captions, entity appearances, and inter-entity relationships, enabling them to consider all relevant aspects within a unified inference step. This consolidated approach aimed to explore the trade-off between a single, comprehensive interaction and the progressive, step-by-step nature of CoA. The specific structure and details of the merged template used for this experiment are illustrated in Fig. 6.
Figure 6: The merged template for single interaction.
Appendix G Additional comparative results
In this section, we present additional comparative results under various settings, including LLaVA-VQA, LLaVA-Caption, Human-labeled, and the proposed CoA. Here, the label set provided by the dataset itself is treated as the Human-labeled labels. The comparative results are illustrated in Fig. 7 and Fig. 8.
For each setting, we compare the predicted labels, calculate the number of correct predictions, and highlight these counts with green numbers. As shown in Fig. 4 and Fig. 5, both LLaVA-VQA and LLaVA-Caption, as well as Human-labeled annotations, tend to focus on predicting prominent features. However, due to the subjectivity of human annotators and the ambiguity in the goal of single-interaction models, these methods result in fewer predicted labels and lower accuracy compared to our proposed CoA. These findings underscore the effectiveness of our approach, which leverages multi-step interactions to achieve more precise and comprehensive semantic label predictions.
Figure 7: Comparison of various settings, including LLaVA-VQA, LLaVA-Caption, Human-labeled and the proposed CoA method. The numbers annotated in the figure represent the count of correctly predicted semantic labels.
Figure 8: Comparison of various settings, including LLaVA-VQA, LLaVA-Caption, Human-labeled and the proposed CoA method. The numbers annotated in the figure represent the count of correctly predicted semantic labels.

### v2

@@ 4.2. Quantitative and Qualitative Results
An interesting observation from the table is that overall, VQA-based approaches perform less effectively than caption-based methods in $\text{M}_{\text{acc}}$. This is primarily because the GSLs task emphasizes identifying all entities relevant to an image. Single-question VQA interactions often lead VLMs to focus on specific local details within an image, whereas captioning guides the model to capture a more comprehensive set of features, resulting in outputs that encompass all relevant entities. This insight influenced the design of our action chain, where the first action is dedicated to obtaining a broad image description. Our experiments validate the effectiveness of this setup.
4.3. Results on Real-world Open-Vocabulary
Table 5 presents a performance comparison on real-world open-vocabulary datasets across two metrics. As can be seen from Table 5, the proposed CoA significantly outperforms all baselines across both metrics. Specifically, the proposed CoA outperform the second-best method by $6.7\%$ ($\text{M}_{\text{com}}$) and $7.73\%$ ($\text{M}_{\text{acc}}$). The notable gains in a real-world scenario highlight CoA’s ability to handle diverse, unconstrained inputs more effectively than existing vision-language models.
Table 4. Generalizability with different VLMs on the VOC datatset.
VLM Vehicles Animals Indoor Outdoor Avg
Mcom $\uparrow$ Macc $\uparrow$ Mcom $\uparrow$ Macc $\uparrow$ Mcom $\uparrow$ Macc $\uparrow$ Mcom $\uparrow$ Macc $\uparrow$ Mcom $\uparrow$ Macc $\uparrow$
LLaVA-v1.5-7B (Liu et al., 2024c) 55.22 66.50 53.16 62.05 54.99 66.66 55.87 66.9 54.81 65.53
LLaVA-v1.5-7B w/ CoA 81.23 (+26.01) 75.18 (+8.68) 79.83 (+26.67) 78.54 (+16.49) 84.51 (+29.52) 79.69 (+13.03) 84.23 (+28.36) 76.54 (+9.64) 82.45 (+27.64) 77.49 (+11.96)
LLaVA-v1.5-13B (Liu et al., 2024c) 68.74 69.67 68.34 67.77 77.95 66.39 68.62 68.5 70.91 68.08
LLaVA-v1.5-13B w/ CoA 81.42 (+12.68) 76.94 (+7.27) 81.63 (+13.29) 79.72 (+11.95) 86.94 (+9.99) 84.35 (+17.96) 90.76 (+22.14) 79.70 (+11.20) 85.19 (+14.28) 80.18 (+12.10)
LLaVA-v1.6-7B (Liu et al., 2024b) 74.00 70.89 71.96 75.32 79.33 70.16 72.63 71.58 74.48 71.98
LLaVA-v1.6-7B w/ CoA 82.15 (+8.15) 76.99 (+6.10) 86.19 (+14.23) 80.66 (+5.34) 86.90 (+7.57) 84.67 (+6.51) 86.79 (+14.16) 75.75 (+4.17) 85.51 (+11.03) 79.52 (+7.54)
Table 5. Comparison of results on real-world open-vocabulary scenario datasets.
Method Macc $\uparrow$ Mcom $\uparrow$
VQA BLIP-2 (Li et al., 2023) 70.76 60.54
InstructBLIP (Dai et al., 2023) 64.40 62.70
LLaVA (Liu et al., 2024c) 71.68 60.33
MiniGPT-4 (Zhu et al., 2024b) 76.76 49.97
Caption BLIP-2 (Li et al., 2023) 72.68 63.01
InstructBLIP (Dai et al., 2023) 72.24 65.73
LLaVA (Liu et al., 2024c) 74.76 61.71
MiniGPT-4 (Zhu et al., 2024b) 77.84 43.62
CoA (Our) 83.30 70.43
Figure 5. Comparison of multiple interaction (CoA) with single merge interaction on the VOC datatset.
Table 6. Computational cost of different methods.
@@ 4.3. Results on Real-world Open-Vocabulary
  VQA   BLIP-2 (Li et al., 2023)   6650 $\pm$ 117
  LLaVA (Liu et al., 2024c)   753 $\pm$ 14
  MiniGPT-4 (Zhu et al., 2024b)   11216 $\pm$ 335
  Caption   BLIP-2 (Li et al., 2023)   6870 $\pm$ 177
  LLaVA (Liu et al., 2024c)   3073 $\pm$ 152
  MiniGPT-4 (Zhu et al., 2024b)   14760 $\pm$ 1433
  Single merge interaction   4826 $\pm$ 561
  LLaVA + CoA (Our)   6110 $\pm$ 149
4.4. Effectiveness Analysis of CoA
To validate the effectiveness of the CoA, we decomposed it into multiple independent components. Since Action 5 serves as the primary step for the GSLs task, responsible for identifying all entity labels in the test image, we used it as the baseline. We then incrementally added each subsequent component to the CoA and evaluated the performance impact after each addition. The results of these evaluations are shown in Table 3. From Table 3, it is evident that each added component contributes to significant improvements in both $\text{M}_{\text{acc}}$ and $\text{M}_{\text{com}}$ metrics. Notably, the complete action chain outperforms using only Action 5, achieving an increase of $13.46\%$ in the $\text{M}_{\text{acc}}$ metric and $2.15\%$ in the $\text{M}_{\text{com}}$ metric. These findings strongly highlight the performance enhancements brought by the CoA, which proves the effectiveness of the proposed method.
4.5. Generalizability with different VLMs
To evaluate the adaptability of CoA across various VLMs, we conducted experiments by changing the base model and testing its performance on different VLMs. As shown in Table 4, all base models (including LLaVA-v1.6 (Liu et al., 2024b)) exhibit significant performance improvements when enhanced with CoA. These findings demonstrate the adaptability and effectiveness of the proposed CoA across diverse VLMs, highlighting its capability to integrate seamlessly with different model architectures.
4.6. CoA vs Single Merge Interaction.
To validate the advantage of multiple interactions, we designed an experiment inspired by the concept of In-Context Learning. This involved merging the multiple actions of CoA into a single interaction (i.e., guiding VLMs to simultaneously consider captions, the appearance of entities, and relationships between entities within a single interaction). The comparison of single-interaction performance versus multi-interaction CoA is reported in Fig. 5. The results clearly demonstrate that multiple interactions provide significant performance benefits over a single-step approach, attributed to the more granular processing and well-defined task objectives that facilitate more accurate outputs from VLMs.
4.7. Computational cost of CoA
Given that the proposed CoA method involves multiple interactions with VLMs, it is essential to evaluate the time cost associated with the overall execution process. We analyze the computational efficiency of CoA versus VQA-based VLMs and Caption-based VLMs and report their inference time in Table 6. Notably, our approach maintains a comparable execution efficiency despite multiple interactions. This is attributed to the simplicity of each individual interaction, ensuring that the total inference time remains within an acceptable range.
5. Conclusion
@@
In this work, unlike traditional VLMs that rely on predefined labels set, we propose a new task, GSLs, which generate comprehensive and contextually relevant labels from an unconstrained semantic space. Besides, we introduced an innovative Chain-of-Action (CoA) method designed to handle the challenging GSLs task. By breaking down the generative task into sequential actions that iteratively extract and refine key information, our method ensures that VLMs generate more accurate and complete semantic labels. Finally, extensive evaluations on widely-used benchmark datasets validate the efficacy of CoA, showcasing significant performance gains across multiple key metrics. This highlights the potential of our method to improve label generation in dynamic, undefined environments.
Limitation and future work. In this work, we do not explicitly consider semantic labels at coarse- and fine-grained. However, in certain application scenarios, coarse- and fine-grained semantic labels can serve complementary roles. Capturing fine-grained semantic distinctions remains a critical challenge in multimodal understanding. Future work may enhance CoA’s granularity and consistency by incorporating explicit user intent signals and memory-aware mechanisms.
6. Acknowledgments
This work was supported by the National Natural Science Founda-tion of China (No. 62306320, 61976217), the Open Project Program of the State Key Laboratory of CAD&CG (No. A2424), and the Natural Science Foundation of Jiangsu Province (No. BK20231063).